### v1

# Educating LLMs like Human Students: Structure-aware Injection of Domain Knowledge

Abstract
This paper presents a pioneering methodology, termed StructTuning, to efficiently transform foundation Large Language Models (LLMs) into domain specialists. It significantly minimizes the training corpus requirement to a mere 0.3% while achieving an impressive 50% of traditional knowledge injection performance. Our method is inspired by the educational processes for human students, particularly how structured domain knowledge from textbooks is absorbed and then applied to tackle real-world challenges through specific exercises. Based on this, we propose a novel two-stage knowledge injection strategy: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT). In the SCPT phase, we organize the training data into an auto-generated taxonomy of domain knowledge, enabling LLMs to effectively memorize textual segments linked to specific expertise within the taxonomy’s architecture. Subsequently, in the SSFT phase, we explicitly prompt models to reveal the underlying knowledge structure in their outputs, leveraging this structured domain insight to address practical problems adeptly. Our ultimate method has undergone extensive evaluations across model architectures and scales, using closed-book question-answering tasks on LongBench and MMedBench datasets. Remarkably, our method matches 50% of the improvement displayed by the state-of-the-art MMedLM2 on MMedBench, but with only 0.3% quantity of the training corpus. This breakthrough showcases the potential to scale up our StructTuning for stronger domain-specific LLMs. Code will be made public soon.
1 Introduction
Large language models (LLMs) have recently seen extensive deployment across various applications [42, 1, 20, 6]. When adapting foundational models (e.g., Llama series [39, 40, 2]) for specialized AI assistants in distinct domains [35, 13], developers usually employ two techniques to enhance LLMs’ proficiency: retrieval-augmented generation (RAG) [25] and domain knowledge injection [14]. While RAG effectively utilizes an external knowledge base to augment information, the retrieval process’s inherent noise poses challenges to generating reliable responses[55, 8]. Thus, another avenue of research focuses on injecting new knowledge into models via training techniques [11, 18, 30].
Continual pre-training [37, 19] has been preferred for integrating new, domain-specific knowledge into existing LLMs [9, 44, 35]. Nevertheless, it often involves costly auto-regressive training on billions of tokens aimlessly curated from the internet rather than from precise domain-specific textbooks [21]. For example, MMedLM [35] curates 25.5B tokens to derive a medical model, and DeepSeek-Coder [13] uses 2T tokens for coding adaptation. The common failure to learn effectively from limited textbook content has been attributed to insufficient data diversity [58], which however violates the observation during the human education process in Fig. 1: students gain knowledge by sequentially studying from textbooks, reviewing knowledge points and structures, and applying this knowledge through proper exercises. In this process, all the new data to learn is textbooks (structured content) and exercising examples (question-answering pairs), and students just adopt their world knowledge to memorize, understand, and apply the knowledge to become domain experts [23, 52].
Figure 1: Discrepancy between human learning process and vanilla LLM adaptation paradigm. Human students learn structured knowledge through textbooks section by section, with particular exercises on related knowledge points. Traditional LLM adaptation involves continual pre-training on data chunks from randomly concatenated text segments, with aimless supervised fine-tuning for conversation alignment. The inherent property of structured knowledge is ignored.
Inspired by this, we propose to inject the domain knowledge from textbooks into LLMs, like educating a human student, through a novel two-stage training strategy: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
In the SCPT stage, we argue that high-quality textbook data alone can adequately infuse initial domain knowledge [12], where the organization of training corpora is crucial. Traditionally, text corpora are merely concatenated and divided into chunks of 2048 [35] or 4096 [13], while the inherent structure of the texts (e.g., catalogs of textbooks) is disregarded. Instead, we propose an automatic approach to maintain each chunk’s knowledge structure. We view each chunk as a knowledge point, employing advanced LLMs to extract domain knowledge taxonomy from the corpus efficiently, bypassing the need for manual annotation. Subsequently, LLMs are trained to predict textual content under the condition of linked knowledge points within the taxonomy, integrating individual training chunks with the entire knowledge framework. Ultimately, models are asked to memorize the knowledge structure to review the whole domain knowledge system.
In the SSFT stage, the goal shifts from knowledge injection to enabling LLMs to recall and utilize their acquired knowledge to tackle real-world challenges. We explicitly elicit knowledge structures in LLMs’ responses, as a beacon for models to targeted information retrieval or logical reasoning for reliable responses. To this end, we derive a scalable strategy to generate question-answer pairs as practice exercises by advanced LLMs such as GPT4 [1] or LLaMA3 [2]. In the scenarios with existing QA pairs like MMedBench [35], we retrieve related knowledge structures and content, instructing LLaMA3 to provide explanations from questions to answers based on these structures. For datasets lacking specific QA samples like LongBench [4], we randomly select knowledge branches from the domain taxonomy and prompt LLaMA3 to craft question-answer-explanation triplets for exercises.
Our ultimate approach, termed StructTuning, outperforms conventional methods in domain knowledge injection by emulating human learning processes through SCPT and SSFT phases. We extensively evaluate StructTuning’s effectiveness across different model architectures and sizes. For domain-adapted language models, we first examine their capability to recall injected knowledge through open-ended QA on the LongBench [4] dataset, then assess their application of this knowledge in addressing real-world issues via multiple-choice QA on the MMedBench dataset [35]. Results from both evaluations underscore the superiority of StructTuning. Remarkably, our strategy achieves a 50% improvement in knowledge injection compared to MMedLM2, the current benchmark in the medical domain, using merely 0.3% of the training data requirement. These findings underscore the potential to scale our method for enhancing domain-specific AI assistants significantly.
Our contribution is summarized as follows:
  • We proposed a novel two-stage training strategy, SCPT and SSFT, to inject domain knowledge into LLMs by preserving and utilizing the inherent structure of the training corpus.
  • We developed a scalable data construction framework to generate structure-aware training samples from original corpora, so as to facilitate the SCPT and SSFT stages.
  • We conducted extensive investigations on our StructTuning strategy on various data and model settings, and achieved competitive knowledge injection performance on MMedBench only using 0.3% of the training corpus.
2 Related Work
Domain Adaptation for Large Language Models. While pre-trained LLMs possess promising capabilities, their performance is often hampered by the scope and recency of their training data, which particularly affects smaller models in downstream applications [56, 43]. Continual Pre-Training (CPT) addresses this by perpetually updating a pre-trained model with domain-specific content [37, 50], with parameter-efficient tuning methods devised to curtail training costs [18, 29]. To keep pace with the latest information, models can be fine-tuned with supervised instruction-response pairs (SFT), thus staying current with the advancing knowledge landscape [30, 35]. Existing literature confirms that combining CPT and SFT is effective for LLMs to remain precise and up-to-date in dynamic fields like law [9, 32], finance [48, 26], medicine [44, 35], and coding [36, 13]. Our study builds upon this CPT-SFT framework, innovating with SCPT-SSFT strategies to efficiently and effectively infuse domain knowledge with the inherent structure hierarchy.
Conditional Language Modeling. The idea of continual pre-training language models on domain corpus in the condition of the knowledge structure is mainly inspired by CTRL [22]. Keskar et al. [22] demonstrates the effectiveness of steering text generation through control codes (one or two words) that signify the desired genre, style, or task. In the era of LLM, system prompt plays a similar role in controlling models’ responses to adapt to different needs and functionalities, such as role-playing, language style transfer, task setting, and behavior setting [7, 45, 3]. Our SCPT approach extends the control codes or system prompts to domain-specific knowledge structures, so as to guide the learning process and tailor the model’s output more closely to specialized fields.
Data Augmentation and Synthesis. Due to the lack of high-quality datasets, data augmentation has emerged as a promising solution to mimic real-world patterns [28]. Traditionally, these methods aim to artificially expand the training dataset size [49, 31] or generate entirely new samples that could help models learn better or adapt to specific tasks [38]. Yet, such methods often overlook the structured nature of domain knowledge, and the aimlessly generated samples may also lack diversity [33, 30], leading to potentially suboptimal training outcomes when applied for domain adaptations [30, 38]. Unlike these traditional approaches, our SSFT design is an innovative departure to address the challenge of retaining and utilizing the structured knowledge inherent in domain-specific content.
3 Methodology
This section presents our methodology to inject domain knowledge into pre-trained LLMs by utilizing the inherent knowledge structure, and Fig. 2 depicts our StructTuning framework. For a limited set of curated domain corpora (typically textbooks), we first develop an automatic approach to extract the knowledge structure, and associate the training chunks to the extracted knowledge structure and points (Sec. 3.1). Then, we design a two-stage training strategy to inject the highly structured domain knowledge into language models by mimicking the human education process, comprising the SCPT (Sec. 3.2) and SSFT (Sec. 3.3) techniques. The details are introduced as follows.
Figure 2: Framework for structure-aware knowledge injection. We extract the inherent knowledge structure in the training corpus, and associate training chunks to corresponding knowledge points. Models are continually pre-trained on data chunks in the condition of the knowledge structure, and fine-tuned with supervised QA samples to elicit their learned knowledge to solve real-world questions (including knowledge-intensive (KI) QA, 2- or multi-hop QA, etc.).
3.1 Automatic Extraction of Knowledge Structure
For web-crawled corpus, the meta-info of knowledge structures (e.g., the table content for a textbook) is usually inaccessible or inapplicable due to the data-parsing quality limitation, and all we have are those sequentially arranged text segments (e.g., page-by-page-chunked content). As shown in Fig. 2 (a), we aim to extract (or, recover) the knowledge structure from the raw corpus for subsequent domain knowledge injection, which is achieved by the following steps.
First, we use the spaCy library [^1] https://github.com/explosion/spaCy to split the content from a textbook at the sentence-level, and merge the sentences to form training chunks within a maximum size (e.g., 2048 tokens [35]). After that, we prompt the advanced Llama3-70B [2] model to summarize the title for each chunk, where the textual content with the abstractive title contributes to a “knowledge point”.
Then, we explore an automatic method to aggregate the sequential knowledge points and extract the inherent structure hierarchy by leveraging advanced language models. Particularly, we take the title list of those text chunks as the input (due to the limitation of context length) to instruct Llama3-70B [2] to identify the inherent structure within the title/chunk list, through in-context learning on the provided examples. The prompt template is displayed in Fig. 3, and the full version is in Appendix D.
Through the top-down structurization on the title list, we manage to extract the hierarchical knowledge structure from the original training corpus. Fig. 3 presents an example of the extracted structure, where we use the tree-like mindmap structure [46] to present the knowledge taxonomy from a textbook. The whole process does not involve human annotation, which reduces the cost and makes our method scalable for handling larger amounts of domain training corpora.
Figure 3: Left: prompt template to extract hierarchical knowledge structures from the given content. Right: example of representing the extracted knowledge structure by using a mindmap template.
After automatically extracting the domain knowledge structure and associating the original training chunks to related knowledge points, we delve into injecting domain knowledge through structure-aware continual pre-training (SCPT) and structure-aware supervised fine-tuning (SSFT).
3.2 Structure-aware Continual Pre-Training
In conventional knowledge injection methods, training corpora are randomly concatenated and chunked into text segments without distinguishing the original content, leading to the fact that models can only absorb domain knowledge that is emergent in the data diversity [33, 30, 35]. In this section, we present another solution to inject knowledge from limited pieces of textbooks by leveraging the highly abstractive and exhaustive domain knowledge structures for continual pre-training.
Figure 4: Example of prompt templates to bridge mindmap structure and textual contents.
We first transform the knowledge structure into natural languages using the same mindmap template in Fig. 3, and prepend it to each training chunk, forcing LLMs to memorize the textual content (knowledge points) in the condition of the associated knowledge path in the structure hierarchy. We collected 20 diversified templates from GPT-4 [1] to bridge mindmap structures and training chunks, one of which is displayed in Fig. 4. The prepended mindmap, as well as the template, does not produce auto-regressive loss. Losses are only calculated in the content part. Formally, we turn the original language modeling in vanilla CPT to conditioned modeling [22] in our SCPT stage:
$p({\bm{x}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i})}\quad\Longrightarrow\quad p({\bm{x}}^{k}|{\bm{s}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i},{\bm{s}}^{k})}$ (1)
where $p({\bm{x}}^{k})$ models the probability distribution for the $k$-th chunk ${\bm{x}}^{k}=(x^{k}_{1},\cdots,x^{k}_{n})$ via the chain rule of probability [5] on each token $x^{k}_{i}$, and ${\bm{s}}^{k}$ denotes the associated knowledge mindmap paths.
As illustrated in Fig. 2 (b), after traversing the $m$ knowledge points in extracted structures, models are asked to recall the whole knowledge hierarchy, i.e., to model the composed probability distribution:
$p(\bar{{\bm{s}}})=\prod_{k=1}^{m}{p({\bm{s}}^{k})}$ (2)
In SCPT, we mimic the human education process to inject knowledge into LLMs in a section-by-section manner, and replay the entire knowledge structure for the models to review and summarize the learned domain knowledge. These two steps iteratively alternate throughout training epochs.
Next, we will introduce how to teach LLMs to explicitly utilize their domain knowledge, which is learned in the SCPT stage, to solve practical problems by doing exercises with our SSFT technique.
3.3 Structure-aware Supervised Fine-Tuning
In traditional knowledge injection paradigms, supervised fine-tuning serves as a tool to align the (continually) pre-trained models to interactive chatbots, e.g., through massive question-answering exercises [9, 35]. However, most QA data augmentation or generation strategies aim at enlarging the quantity and enhancing the diversity of training samples [49, 31, 28], which neglects the nature of the highly structured domain knowledge. Therefore, our structure-aware supervised fine-tuning (SSFT) technique focuses on eliciting models’ structured knowledge learned during the SCPT stage, adapting LLMs to interactive and reliable domain experts.
Fig. 2 (c) illustrates the idea of synthesizing SSFT samples guided by domain knowledge structures, as extracted in Sec. 3.1. Specifically, we use the random walk algorithm to create knowledge paths with $1$ to $l$ branches in the original mindmap. For paths with only one knowledge point, we use the corresponding text content to prompt Llama3-70B [2] to generate knowledge-intensive question-answering pairs. For paths with two or more branches, we prompt Llama3-70B with the knowledge path and textual contents to synthesize 2- or multi-hop QA samples, which require specific reasoning along the knowledge structure to derive from questions to answers. Fig. 5 presents several examples.
Figure 5: QA samples synthesized for SSFT. We instruct Llama3-70B to generate (a) knowledge-intensive and (b) multi-hop questions and derive the diagnosis answers with explicit reasoning.
For every synthesized QA sample (${\bm{z}}$), we will prepend the relevant mindmap hierarchy to the answer, and add a CoT prompt in the question to construct another type of QA data (${\bm{z}}^{\prime}$) for SFT alignment. This design explicitly elicits the learned knowledge in models’ responses, teaching them how to apply the structured knowledge to address real-world problems. We use the two types of QA samples for training, as recommended by Qiu et al. [35]. During testing, we can either use the vanilla question as input to efficiently gather models’ answers, or take the CoT prompt to probe to what extent LLMs can memorize and leverage the injected knowledge to answer the questions.
Integrating with SCPT and SSFT, our StructTuning approach translates into remarkable efficacy and efficiency in domain knowledge injection, as comprehensively evaluated in the following sections.
4 Experiments
We design a comprehensive evaluation of our StructTuning through seveal experiments on two benchmarks. First, we investigate the free-form question-answering task on the LongBench [4] dataset, in order to verify the memorization and understanding of injected knowledge (the answer can be directly found in training corpora). Then, we delve into the multi-choice question-answering task on MMedBench [35], to explore how LLMs apply the injected knowledge in basic medicine to determine the real-world diagnosis for patients with logical reasoning.
4.1 Preliminary Investigation on Free-form Question-Answering
Datasets and Tasks. LongBench [4] is a multi-task benchmark tailored for open-book reading comprehension evaluation, where LLMs generate answers to given questions based on one or several lengthy passages. To focus on knowledge injection, we turn the open-book evaluation into a closed-book QA task, where contextual passages are injected into LLMs with training techniques, and models answer questions without input passages. We choose 7 subsets with 1,350 test examples from LongBench for single- and multi-document question-answering evaluation, and the remaining synthetic or code-orientated tasks are eliminated. More details are described in Sec. A.1.
Evaluation Metrics. For free-form QA in LongBench [4], we report the F1-Score between models’ outputs and ground-truth answers for both open-book and closed-book QA tasks. Besides, we also report the Recall for model responses to quantify the memorization degree of injected knowledge.
Investigated Models. For LongBench, we mainly investigate the knowledge injection to the Llama2-7B model [40], and report the Open- and Closed-Book QA performance for comparison.
Implementation Details. On LongBench for closed-book QA, we continually pre-train the Llama2-7B model on 10,476 passages for 3 epochs using a batch size of 128. The initial learning rate is 2e-5, decayed to 0 at the end of training with a cosine scheduler. Then, we query Llama3-70B to generate 2,700 SSFT QA examples, and eliminate the generated data whose answer shares over 0.5 F1-Score with one of the test samples to avoid data leakage. We train the knowledge-injected models for 1 epoch to avoid overfitting. Models are trained with the LLaMA-Factory framework  [^2] https://github.com/hiyouga/LLaMA-Factory.
Figure 6: An example for structure-aware responses.
Main Results. In Tab. 1, we first report the Open-Book QA (OBQA) baseline for Llama2-7B with passages inputs, which shows the best performance on MultiFieldQA (MFQA) [4] and 2WikiMultihopQA (2Wiki) [17] subsets. Then, we establish the Closed-Book QA (CBQA) baseline for pre-trained Llama2-7B, which aims to identify the knowledge that has already been equipped from the newly injected knowledge. Consequently, we follow the traditional CPT+SFT paradigm to inject passage contents into model parameters, and supplement the experiment of our SCPT+SSFT technique for comparison. According to the results shown in Tab. 1, the traditional continual pre-training strategy immediately improves the QA performance on six out of seven subsets (except the Qasper [10] subset), while the supervised fine-tuning alignment does not bring significant gains on the overall F1-Score measure. In contrast, our SCPT+SSFT approach successfully boosts the closed-book QA performance to 20.1% on average, even surpassing the open-book QA baseline of 18.7%. It indicates the vanilla SFT strategy can only regularize LLMs’ response styles, while our SSFT could teach LLMs to utilize their knowledge (injected in the SCPT stage) to answer corresponding questions, as exemplified in Fig. 6.
Table 1: F1 Score evaluation of Open-Book QA (OBQA) and Closed-Book QA (CBQA) tasks on the LongBench [4] dataset. The best results are marked in bold, and the secondary results are marked with underlines. The backbone model is Llama2-7B [40].
Task Adaptation SingleDoc-QA MultiDoc-QA Average
Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
OBQA - 19.2 36.8 11.9 25.4 32.8 9.4 5.2 18.7
CBQA - 17.3 15.9 1.4 5.7 6.7 3.3 1.8 7.4
CPT 16.8 23.6 17.5 18.0 18.6 6.7 14.3 16.5
CPT+SFT 16.8 23.1 13.2 21.3 19.1 10.4 13.4 16.8
SCPT 15.2 21.5 15.2 14.9 19.8 5.6 14.1 15.2
SCPT+SSFT 19.7 23.5 19.5 26.4 24.1 12.2 15.4 20.1
Table 2: Recall evaluation of the Closed-Book QA task. The best results are marked in bold.
Adaptation SingleDoc-QA MultiDoc-QA Average
Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
CPT 13.4 33.0 16.6 28.1 27.7 15.9 13.1 21.1
CPT+SFT 20.7 35.3 20.6 29.9 32.1 18.9 12.0 24.2
SCPT 18.8 42.5 17.7 35.7 36.4 20.5 15.3 26.7
SCPT+SSFT 30.5 44.6 24.3 40.8 42.0 21.8 16.8 31.5
To further support this claim, we compute Recall between model outputs and ground-truth answers in Tab. 2 as a proxy measure for memorizing and manipulating the injected knowledge [59]. Specifically, we use a CoT instruction of Answer the question and explain why. to elicit models’ knowledge in their responses. As shown in Tab. 2, our SCPT strategy has already achieved higher knowledge recall than the CPT-SFT paradigm (26.7% v.s. 24.2%). It implies our structure-aware continually pre-trained model has successfully associated the relevant passages with their entire knowledge structure for the given question, which provides the knowledge path to seek targeted information to derive the answer. As displayed in Fig. 6, our SSFT technique explicitly teaches the model to recall the learned knowledge and answer the questions, which further improves the knowledge recall in Tab. 2.
Figure 7: MindMap Recall
F1-Score BERTScore
0.61 0.87
In addition, we also statistically quantify the memorization of injected knowledge structures by comparing the mindmap in models’ responses with the ground-truth answers via lexical ROUGE-L [27] and semantic BERTScore [54] measures. The results shown in Fig. 7 indicate a relatively good memorization of the injected knowledge mindmap, showing the efficacy of our SCPT strategy.
4.2 In-depth Evaluation for Multi-choice QA Application
Datasets and Tasks. MMedBench [35] is a multilingual medical multi-choice QA benchmark, with 45,048 QA pairs for adapting LLMs to medical experts and 8,518 for testing. We collect 76M textbook corpora from MedTextBooks [21] and MMedC [35] for medical knowledge injection and use the training/test split from MMedBench for SFT/evaluation. Detailed setup is in Sec. A.2.
Evaluation Metrics. For the multi-choice QA task in MMedBench [35], we follow the default setting to calculate the accuracy on six language subsets, as well as the averaged scores.
Investigated Models. For MMedBench, we extend the investigated LLMs across model scales and architectures including Llama2-7B, Llama2-13B [40], and InternLM2-7B [57]. We also compare our knowledge-injected models with other open- or closed-source LLMs, such as the domain-specified MedAlpaca [15], ChatDoctor [53], PMC-LLaMA [47], and MMedLM [35] models.
Implementation Details. We first train LLMs for 3 epochs on medical textbooks with a batch size of 128. Then, we ask Llama3-70B to create structure-aware explanations for existing 45K QA samples in MMedBench’s training split and 33K extra entries by traversing extracted knowledge structures. Syntheses with overlapped options in the test set are removed. In the SFT phase, the learning rate is set as 1e-6 to avoid overfitting on such an amount of SFT samples, as suggested by Qiu et al. [35].
Main Results. In Tab. 3, we present the overall performance across a series of LLMs on the testing split of MMedBench. The results demonstrate the promising enhancement achieved by our StructTuning technique, which translates into consistently significant improvements on Llama2-7B (+8.78%), Llama2-13B (+6.17%), and InternLM2-7B (+4.46%). It shows the generalizability and scalability of our StructTuning strategy across model architectures and sizes. Notably, our method, comprising SCPT on 76M curated textbook corpora and SSFT on 78K QA samples, achieves 50% performance (4.46% v.s. 8.76%) against the state-of-the-art MMedLM2-7B model, which is trained on a huge MMedC [35] corpora of 25.5B tokens. Our approach shows much more efficiency in transforming pre-trained LLMs into domain experts by structured-aware knowledge injection, which largely reduces the cost for LLMs’ domain adaptation to 0.3%. However, our method is currently unable to achieve 100% of knowledge injection efficacy against MMedLM2, which may comprise two major reasons: (1) some of the knowledge in the training corpus has already been learned by the foundation InternLM2 model, and (2) the knowledge in our curated 76M training data cannot cover all the testing samples. We leave the promotion of achieving 100% improvement in our future work.
Table 3: Multiple-choice accuracy evaluation on MMedBench [35]. We report each model’s accuracy across six languages separately, with “Average” denoting the mean score over six languages. We also list out the data quantity for CPT and SFT stages during medical knowledge injection.
Model CPT SFT English Chinese Japanese French Russian Spanish Average
InternLM-7B - 45K 44.07 64.62 37.19 24.92 58.20 44.97 45.67 +0.00
MMedLM-7B 25B 45K 49.88 70.49 46.23 36.66 72.27 54.52 55.01 +9.34
InternLM2-7B - 45K 57.27 77.55 47.74 41.00 68.36 59.59 58.59 +0.00
MMedLM2-7B 25B 45K 61.74 80.01 61.81 52.09 80.47 67.65 67.30 +8.71
Llama2-7B - 45K 43.36 50.29 25.13 20.90 66.80 47.10 42.26 +0.00
+Ours 76M 78K 49.41 65.15 36.68 35.21 69.14 50.62 51.04 +8.78
Llama2-13B - 45K 51.37 57.97 32.66 25.08 69.92 52.99 48.33 +0.00
+Ours 76M 78K 53.02 68.30 37.78 41.71 70.70 55.51 54.50 +6.17
InternLM2-7B - 45K 57.27 77.55 47.74 41.00 68.36 59.59 58.59 +0.00
+Ours 76M 78K 60.80 79.19 50.75 45.34 75.39 66.85 63.05 +4.46
Ablation Studies. We conduct a comprehensive ablation study with the English split of the MMedBench dataset to evaluate our efficacy. Specifically, we take Llama2-7B as the backbone model, select the English textbooks [21] (with 26M tokens) for vanilla and our structure-aware continual pre-training, and use different QA samples for supervised fine-tuning. In particular, in Tab. 4, SFT refers to vanilla SFT with 10K English QA samples provided in MMedBench’s [35] training split, SSFT indicates structure-aware SFT on 10K training data, where the questions are the same as SFT while the answers are enhanced with knowledge structure explanation by Llama3-70B, as described in Sec. 3.3. SSFT* includes another 8K structure-aware QA syntheses by Llama3-70B, consisting of 18K entries for training. The training hyper-parameters follow the main experiment.
Results in Tab. 4 imply the vanilla CPT+SFT injection paradigm does bring considerable improvement of 1.73% in the English split, while our SCPT technique with vanilla SFT strategy presents a slightly higher accuracy of 46.50%. However, when combining SSFT with SCPT, our method immediately leads to a significant boost in the English split (49.96% v.s. 44.54%), demonstrating the efficacy and necessity of eliciting the learned structured knowledge to solve practical problems. Moreover, the supplementation of 8K extra QA syntheses surprisingly enhances the model performance on the other five subsets, resulting in a valuable gain on the average accuracy of 38.27%. This result demonstrates the knowledge transferability across different languages [24, 34]. After training with our SSFT strategy, LLMs can actively utilize the knowledge injected in one language to solve the problem in another language, further evidencing our superiority against the traditional SFT technique.
In addition, we observe that the commonly used RAG [25] strategy does not bring significant advantages to the MMedBench evaluation. The main reason lies in the gap between the pre-training corpus (comprising official knowledge statements from textbooks) and evaluated QA samples (originating from practical diagnosis records). To teach LLMs to apply relevant medical knowledge for reliable diagnosis answers, knowledge injection by CPT and SFT shows more advantages in this situation.
Table 4: Ablation studies with Llama2-7B on the English subset of MMedBench. The best results are marked in bold, and the secondary results are marked with underlines.
Adaptation English Chinese Japanese French Russian Spanish Average
- SFT 44.54 32.81 26.63 15.27 53.91 42.30 35.91
CPT SFT 46.27 32.57 26.13 17.36 50.00 40.63 35.49
SCPT SFT 46.50 32.14 20.10 18.17 53.91 39.97 35.13
SCPT SSFT 49.96 32.63 22.11 17.52 51.17 41.28 35.78
SCPT SSFT* 49.10 33.92 18.33 27.14 57.42 43.73 38.27
RAG 38.12 29.22 22.61 23.34 53.91 36.47 33.95
5 Conclusion
This work pioneers in incorporating structure-aware methodologies to enhance domain knowledge injection into large language models (LLMs). Through a novel two-stage training strategy, combining the SCPT and SSFT techniques, we have set a new precedent for efficiently adapting LLMs to specialized domains. The promising results and the scalable data construction framework we developed underscore the viability and potential of our method. The limitations and broader impacts are discussed in Appendix C. We hope to inspire further research in efficient and effective domain adaptation, moving a step closer to models that can truly emulate human intelligence.
Appendix A Implementation Details
A.1 Detailed Setup on LongBench
To focus on the investigation of knowledge injection, we choose 7 subsets from LongBench [4] across single- and multi-document QA tasks in English and Chinese, and the remaining synthetic or code-orientated tasks are eliminated:
  • Single-Doc QA. For single-document QA, we take three subsets from LongBench: (1) Qasper [10], featured by question-answering over NLP technical papers and annotated by NLP practitioners; (2) MultiFieldQA [4], manually curated from multiple data sources and annotated by Ph.D. students; and (3) MultiFieldQA-zh, the Chinese split also provided by Bai et al. [4], covering multiple Chinese scenarios. MultiFieldQA contains 150 Context-Question-Answer triplets to test, and the others adopted subsets include 200 pieces of test samples respectively.
  • Multi-Doc QA. Multi-document QA requires LLMs to extract and combine information from multiple documents to derive the answer, which is generally more challenging than single-doc QA. We take four multi-hop QA datasets: (1) HotpotQA [51], containing 2-hop questions written by native speakers given two related paragraphs; (2) 2WikiMultihopQA [17], involving up to 5-hop questions synthesized through manually designed templates on Wikipedia passages; (3) MuSiQue [41], carefully composed with up to 4-hop reasoning on an increased number of supporting and distracting context evidence; and (4) Dureader [16], developed based on Baidu Search and Baidu Zhidao and filtered by Bai et al. [4] to reduce the data noise. Each subset has 200 test samples.
A.2 Detailed Setup on MMedBench
Data for Evaluation. The Multilingual Medical Benchmark (MMedBench) [35] represents a comprehensive and diverse multilingual medical Question and Answering (QA) benchmark designed to evaluate models’ capabilities of understanding and processing medical content.
MMedBench’s robust dataset extends across 6 languages (i.e., English, Chinese, Japanese, French, Russian, and Spanish) and 21 medical fields, which include, but are not limited to, Internal Medicine, Biochemistry, Pharmacology, Psychiatry, and many others. It provides 45,048 training pairs and 8,518 testing pairs for diverse learning and testing scenarios. The training split is specifically designed for domain-specific finetuning of large language models (LLMs), while the entire testing set allows for a precise assessment of multi-choice question-answering performance. Statistics on six languages are displayed in Tab. A1. Notably, the benchmark includes scenarios where questions may have multiple correct answers (i.e., in Japanese and French subsets), introducing additional complexity for the model evaluation process.
Table A1: Sample statistics on MMedBench.
Split English Chinese Japanese French Russian Spanish Total
Train 10,178 27,400 1,590 2,171 1,052 2,657 45,048
Test 1,273 3,426 199 622 256 2,742 8,518
Data for Continual Pre-Training. To investigate high-quality domain knowledge injection for LLMs, we collect 18 English textbooks and 33 Chinese textbooks from the National Medical Board Examination in the USA and Mainland China, respectively [21]. All collected textbooks are originally in PDF format and Jin et al. [21] converted them into digital text via OCR and performed some clean-up pre-processing strategies to reduce the data noise. The English and Chinese textbooks count for around 26.1M and 21.5M tokens by Llama2 tokenizer [40]. Then, we randomly sample an extra 28M textbook corpora from MMedC [35] for the other languages (except the unavailable Japanese textbooks), and the final statistics are displayed in Tab. A2.
Table A2: Sample statistics on training data.
Stage English Chinese Japanese French Russian Spanish Total
CPT 26.1M 21.5M - 8.1M 10.3M 10.1M 76.1M
SFT 18.8K 39.1K 1.6K 5.3K 5.9K 7.5K 78.2K
Data for Supervised Fine-Tuning. As introduced in Sec. 3.3, we prompt Llama3-70B [2] to build the structure-aware answer explanations on top of the raw SFT samples in MMedBench’s training split, and generate extra QA pairs by traversing the extracted knowledge structure from textbooks. The final quantity statistics are presented in Tab. A2.
Appendix B Resource Requirement
We use 8 NVIDIA A100-80G GPUs to train all the models, and leverage 1-2 NVIDIA A100-80G GPUs for inference.
Appendix C Limitations and Discussions
Limitations. A noteworthy limitation of our work, despite its significant achievements, is centered around the exploration of the full potential of our StructTuning strategy for knowledge injection. While our method has successfully demonstrated a remarkable 50% improvement in domain-specific knowledge application and retention using merely 0.3% of the training corpus, the extent of data scalability needed to achieve 100% of the desired effectiveness remains unexplored. This limitation primarily arises from constraints in time and resources available during our research phase. We will delve into the investigations in our future work.
Broader Impacts. We discuss the positive and negative societal impacts as follows:
  • Positive Societal Impacts. Our StructTuning strategy could democratize access to advanced AI, enabling specialized AI assistants across various sectors such as healthcare and education. This method’s efficiency could improve diagnostic tools, personalized learning experiences, and enhanced decision-making processes. Furthermore, by mirroring the human learning process, our approach facilitates the development of AI systems that are easier for people to trust and interact with.
  • Negative Societal Impacts. The enhanced capabilities of LLMs might be exploited for generating convincing yet fraudulent content, raising concerns over misinformation and intellectual deceit. Additionally, if the structured knowledge base contains biases, it could inadvertently reinforce societal stereotypes, necessitating rigorous curation to ensure fairness and accuracy. To counteract potential misuse, controlled model releases and continuous monitoring for abuse are recommended. Ensuring the diversity and contemporaneity of training data can mitigate bias. Collaborative efforts with domain experts, ethicists, and policymakers are vital in developing guidelines for the technology’s ethical use.
Appendix D Prompt Template for Structurization
Fig. A1 displays the few-shot prompt template to query giant commercial LLMs to execute structurization. We first introduce the instruction for structurization (including the task definition and output formats), then provide two representative examples to further illustrate the process. The first example in Fig. A2 tells the model to focus on the existing numerical or enumeration indicators to assist in constructing the aspect-description structure. The second example in Fig. A3 teaches the model to automatically summarize the aspects and attach their corresponding descriptions from the raw text sequences. Finally, in Fig. A1 we ask commercial LLMs to structurize the $input\_statement$ from user input as expected.
Figure A1: Prompt template for few-shot structurization.
Figure A2: The first example for few-shot structurization.
Figure A3: The second example for few-shot structurization.